Find the Wronskian of a given set {y1, y2} of solutions of
(1 − x^2)y" − 2xy' + α(α + 1)y = 0 when W(0) = 1
1
Expert's answer
2020-08-06T14:03:43-0400

Solution.

(1-x2)y'' - 2xy' + α(α + 1)y = 0


Dividing both sides of the equation by (1-x2) we get differential equation in standard form:


"y''-\frac{2x}{1-x^2}y'+\frac{\alpha \left(\alpha +1\right)}{1-x^2}y=0" ;


p(t) is the coefficient of y' and q(t) is the coefficient of y:

"p\left(t\right)=-\frac{2x}{1-x^2}"


"q\left(t\right)=\frac{\alpha \left(\alpha +1\right)}{1-x^2}"

To determine the Wronskian we apply Abel's theorem:

"W\left(y_1,y_2\right)\left(x\right)=C\cdot e^{\left(-\int p\left(x\right)dx\right)}"


"W\left(y_1,y_2\right)=C\cdot e^{\left(-\int \left(-\frac{2x}{1-x^2}dx\right)\right)}"


"W\left(y_1,y_2\right)=C\cdot e^{\left(-\int \left(\frac{d\left(1-x^2\right)}{1-x^2}\right)\right)}"


"W\left(y_1,y_2\right)=C\cdot e^{-\ln \left|1-x^2\right|}"


"W\left(y_1,y_2\right)=\frac{C}{\left|1-x^2\right|}"


when W(0) = 1, we will get:

"1=\frac{C}{\left|1-0\right|}"

C = 1, thus:


"W\left(y_1,y_2\right)=\frac{1}{\left|1-x^2\right|}"


Answer: W = 1 / |1 - x2|